Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 100 miles per hour and Train B is traveling at 104 miles per hour. Train A passes a station at 11:10 am. If train B passes the same station at 11:40 am at what time will train B catch up to train A
_____am or _______pm


Can some one please help me. Not sure on how to set up the equation

Answer by checkley77(12844)

D=RT
BECAUSE THE DISTANCES WILL BE EQUAL WE HAVE:
D=100T FOR TRAIN A.
D=104(T-.5) FOR TRAIN B.
100T=104(T-.5)
100T=104T-52
100T-104T=-52
-4T=-52
T=-52/-4
T=13 HOURS OR @ 12:10 AM. THE NEXT MORNING THEY WILL MEET.
PROOF:
100*13=104(13-.5)
1,300=104*12.5
1,300=1,300 MILES TRAVELED.